OpenStudy (anonymous):

x2-x-1=0, this problem is squared

OpenStudy (saifoo.khan):

this is quadratic.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer to this is called the golden ratio

OpenStudy (anonymous):

how do i solve it

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

first imagine the equation as ax^2+bx+c=0 then use the formula x=(-b±√(b^2-4ac))/2a

OpenStudy (anonymous):

the answer is x=1/2(1-sqroot5) x=and1/2(1+sqroot5)

OpenStudy (anonymous):

the answer is an important number in maths called the golden ratio

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2-x-1%3D0

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i have 3 more problems, if you want to help me

OpenStudy (anonymous):

if i can

OpenStudy (anonymous):

okay, 2/3x2(squared) + 1/4x = 3

OpenStudy (anonymous):

i think x=9

OpenStudy (anonymous):

-9 sorry